OpenStudy (anonymous):

Dillon can buy a fixed-rate bond with a 4% coupon, a $10,000 principal, and five years left to maturity or a TIPS with the same principal and age to maturity with a 2% coupon. Compare the two bonds if inflation runs 3% per year on average. Select the best answer from the choices provided. The fixed-rate bond has a better yield. The TIPS has a better yield. Both yields are roughly equivalent. The yields vary with inflation.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 its D ?

OpenStudy (tkhunny):

Sounds like a job for a present value. 10000*0.04/2 = 10000*0.02 = 200 -- Okay, there's the one coupon. Go!

OpenStudy (anonymous):

isn't the answer TIPS? because it adjusts with inflation, making the fixed rate bond to gradually decrease in yield. ( interest rate of fixed bond is 4% - inflation rate of 3% = 1% real rate of interest for fixed rate bond) VS. 3% added to the TIPS bond principal...(10,000 * .03= 300.00 added in principal every year as well as interest rate of 2%) resulting to (TIPS) 10,300 * .02=206 . VS (fixed rate)10,000 * 01=100